OpenStudy (anonymous):

How do you write an equation in slope intercept form for a line that passes through a certain set of points? And how do you do it if they tell you the x intercept, but not the y intercept? Thanks!

OpenStudy (anonymous):

slope =(y2-y1)/(x2-x1) slope intercept form = y=mx+b

OpenStudy (anonymous):

Ok, so you use the points in slope formula? But what if you're only given one set of points? Like if it said, slope -5, passes through (-3,-8) ? then how would you do the formula?

OpenStudy (anonymous):

use formula, \[y-y_{1}=m(x-_x{1})\]

OpenStudy (anonymous):

m=-5 y1=-8 x1=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

u need help with equation?

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

Thanks! ok how do I give you a medal haha?

OpenStudy (anonymous):

next my name click on good answer

OpenStudy (anonymous):

Oh wait, what if it's perpendicular?

OpenStudy (anonymous):

2 line intersect have angle 90 degree

OpenStudy (anonymous):

paralell never intersect

OpenStudy (anonymous):

Right, I know, so how would you put that in slope intercet equation?

OpenStudy (anonymous):

slope -5, passes through (-3,-8) ?

OpenStudy (anonymous):

No, if it says that it's perendicular to lik efor examle, y=3x-2

OpenStudy (anonymous):

let do this ffrist y-(-8)=-5(x-(-3)) y+8= -5(x+3) y= -5x-15-8 y= -5x -23

OpenStudy (anonymous):

Yes, I got that

OpenStudy (anonymous):

u want write anothe equation is perendicular to this line y=3x-2?

OpenStudy (anonymous):

yes, and passes through (6,-1)

OpenStudy (anonymous):

perendicular slope is flip and opposite side Example : slope = -1/3 than y-(-1)=-1/3(x-6) \[y+1=\frac{-1}{3}( X)+2\]\[y=\frac{11x}{3}+2-1 \]\[y=\frac{-1x}{3}+1\] \[y=\frac{1x}{3}+1 \]perendicular to line y=3x-2 u can check by : produc of 2 slope is -1 then they are perendicular

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[\frac{-1}{3}\times3=-1\] -1 yes they are perendicular

OpenStudy (anonymous):

good cluck your home work

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

u forgot give me a medal

OpenStudy (anonymous):

How do I do that again?

OpenStudy (anonymous):

next my name click on good answer

OpenStudy (anonymous):

d u see?

OpenStudy (anonymous):

next may name right hand side

OpenStudy (anonymous):

It won't let me click on the medal symbolk though

OpenStudy (anonymous):

u see medal or good answer?

OpenStudy (anonymous):

click on give medal or good answer

OpenStudy (anonymous):

There is nothing that says that

OpenStudy (anonymous):

not thing next to my name? some one just give me

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

There is a medal symbol and it says 117, and then a shield and it says 10. I can't click on either of those. There is nothing else.

OpenStudy (anonymous):

next it says 10

OpenStudy (anonymous):

after it says 10

OpenStudy (anonymous):

this u see the box say good answe or give medal?